OpenStudy (anonymous):

if y varies directly as x and y=49 when x=7, what is the value of y when x=3? a. 7 b. 9 c. 49 d. 21

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\color{orangered}{\rm k} x}\) this above, is the equation of a direct variation. In this equation x and y are directly proportional to y, so that you can multiply x by some constant k and obtain y. (this perfect proportion is called direct variation) it is a line with a slope k that goes through the origin.

OpenStudy (solomonzelman):

If you are given that that `y=49 when x=7` then you can go ahead and find the constant k. plug in 49 for y and 7 for k, into the \(\large\color{black}{ \displaystyle y=\color{orangered}{\rm k} x}\)

OpenStudy (solomonzelman):

and then after plugging in solve for k

OpenStudy (solomonzelman):

When you find the k, you can figure out the value of y when x=3. How? you can plug in the value of k you just found, and plug 3 instead of x. (again u r plugging into \(\large\color{black}{ \displaystyle y=\color{orangered}{\rm k} x}\) )